# Unsupervised detection of semantic correlations in big data

### v1

@@ Acknowledgements
Code and data availability
Supplementary Information
Optimization
@@
Spin systems
@@
The BID of image representations
We followed the experimental setup of Ansuini et al. [2019] by computing separately the BID of Resnet18 representations from 8 highly populated classes of ImageNet (’vizsla’, ’koala’, ’Shih-Tzu’, ’Rhodesian_ridgeback’, ’English_setter’, ’cabbage_butterfly’, ’Yorkshire_terrier’), where for each class we have roughly 1300 images. We measure the BID after each Resnet block, i.e., after each skip connection. Resnets have ReLU activation functions, so only two possible outputs of each unit are possible: zero or greater than zero. Taking the sign of such an output automatically constructs the binary system under the convention $sign(0)=0$. The preprocessing of images was done following PyTorch official documentationPyt [a]. The layer indexation was taken following the list of "graph_node_names", from the official PyTorch documentationPyt [b].
The BID of text representations
@@
Maximum Likelihood Estimation of a scale-dependent BID
@@
such that $\langle n_{A}\rangle=\rho V_{A}$. We consider a data point $i\in A$, and another set $B$ that contain $A$. Then, if the density is constant in set $B$ ($\rho_{A}=\rho_{B}$ = $\rho$) the conditional probability of having $n_{A}$ points in $A$ given that there are $n_{B}$ points in $B$ is
@@
$\mathcal{L}(n_{i}|k_{i},p_{i})=\prod_{i=1}^{N_{s}}Binomial(n_{i}|k_{i},p_{i}),$ (10)
where $N_{s}$ is the number of samples in the dataset. This equation is maximized to obtain the Maximum Likelihood Estimator $\hat{d}$. Doing this, in the limit of large number of samples, the equation that defines $\hat{d}$ takes the form
@@
Neighbourhoods are approximately preserved after binarizing high dimensional activations

### v2

@@ Acknowledgements
Author contributions
SA performed the numerical experiments. SA, AR, and AL designed the experiments, analyzed the data, and wrote the manuscript.
Competing interests
The authors declare no competing interests.
Appendix A Supplementary Information
A.1 Optimization
@@
A.2 Dependence on the number of samples
Fig. 6 shows the dependence of the BID on the number of samples, $N_{s}$, for the 2D ferromagnetic Ising model. The algorithm provides a stable estimate after a few hundred samples even for high ID values because the fit of the model is performed on a global observable, the probability of observing Hamming distance r in the dataset, $P_{emp}(r)$, a quantity which is well defined even in the presence of measure concentration in high dimensions. The right panel of Fig. 6 shows that the smallest sampled scales are those strongly affected by the size of the dataset, but larger values of $r$ are stable, i.e., the histograms computed for the different values of $N_{s}$ overlap. Note that Eq. (4) weights the contribution of the discrepancies with $P_{emp}(r)$, and thus the most statistically relevant distances dominate the optimization.
Figure 6: Left panel: BID normalized per bit versus the number of samples, for different temperatures, $T$. Right panel: Example histograms for $T=3$, and three different values of $N_{s}$. In solid black lines, the fits to each histogram (they overlap).
A.3 Bayesian inference of the BID by Markov Chain Monte Carlo
Figure 7: Empirical posterior density for the two parameters of our model. $d_{0}^{*}/N=0.0512$ and $d_{1}^{*}=1.9079$ correspond to the solution found optimizing Eq. (2). The posterior means and standard deviations are $\mu_{d_{0}}/N=0.0510$, $\sigma_{d_{0}}/N=0.0006$, $\mu_{d_{1}}=1.908$, $\sigma_{d_{1}}=0.002$ Number of samples: $N_{s}=500$, number of spins $N=10^{4}$, temperature $T=2.3.$
The parameters of model (2) can be found minimizing Eq. (4), as we did in our work, or alternative using the Metropolis Hastings algorithm on model (2), with a flat prior. Here we sample the probability distribution
$P(d_{0},d_{1}|P_{emp}(r))=\prod_{r\in\text{data}}\bigg{(}\frac{\binom{d(r)}{r}}{2^{d(r)}}\bigg{)}^{N_{r}},$ (5)
where $N_{r}$ is the number of times the distances $r$ was sampled in the dataset. Fig. 7 shows the equivalence of the two methods on the 2D ferromagnetic Ising model at $T=2.3$. The MCMC simulation was performed starting at $(d_{0},d_{1})=(\langle r\rangle_{P_{emp}},1)$, making random moves with a maximum amplitude given by 1 percent of the current value of $(d_{0},d_{1})$ at each time step $t$. This approach allows estimating the mean, variance and covariance of the parameters of the model.
A.4 Spin systems
@@ A.4.2 Potts model
A.5 The BID of image representations
We followed the experimental setup of Ansuini et al. [2019] by computing separately the BID of Resnet18 representations from 8 highly populated classes of ImageNet (’vizsla’, ’koala’, ’Shih-Tzu’, ’Rhodesian_ridgeback’, ’English_setter’, ’cabbage_butterfly’, ’Yorkshire_terrier’), where for each class we have roughly 1300 images. We measure the BID after each Resnet block, i.e., after each skip connection. Resnets have ReLU activation functions, so only two possible outputs of each unit are possible: zero or greater than zero. Taking the sign of such an output automatically constructs the binary system under the convention $sign(0)=0$. The preprocessing of images was done following PyTorch official documentation.The layer indexation was taken following the list of "graph_node_names", from the official PyTorch documentation.
A.6 The BID of text representations
@@
A.7 Maximum Likelihood Estimation of a scale-dependent BID
@@
such that $\langle n_{A}\rangle=\rho V_{A}$. We consider a data point $i\in A$, and another set $B$ that contains $A$. Then, if the density is constant in set $B$ ($\rho_{A}=\rho_{B}$ = $\rho$) the conditional probability of having $n_{A}$ points in $A$ given that there are $n_{B}$ points in $B$ is
@@
$\mathcal{L}(n_{A}^{i}|n_{B}^{i},p)=\prod_{i=1}^{N_{s}}Binomial(n_{A}^{i}|n_{B}^{i},p),$ (11)
where $N_{s}$ is the number of samples in the dataset, and identical volumes $V_{A}$ and $V_{B}$ where taken around every data point $i$. This equation is maximized to obtain the Maximum Likelihood Estimator $\hat{d}$. Doing this, in the limit of large number of samples, the equation that defines $\hat{d}$ takes the form
@@
A.8 Comparison with other methods
A.8.1 ID dependence with crop size in image classification
Fig. 11 shows the intrinsic dimension computed using GRIDE and FCI on the same effective spin system of Fig. 2 for an example relative depth of 0.39. For the GRIDE estimator we used the parameter $k=4$ to select the scale.
Figure 11: BID results from Fig 2 compared with FCI and GRIDE on the same spin dataset.
A.8.2 ID of Text representations
Fig. 12 shows the ID estimated by BID,FCI and GRIDE on the effective spin system of Fig. 3. The two curves computed at layer zero (embedding layer) with FCI are cut at 200 tokens since the implementation present in https://github.com/vittorioerba/pyFCI crashed due to overflow.
Figure 12: BID results from Fig 3 compared with FCI and GRIDE on the same spin dataset. For the last layer we have that $BID(T)/N\propto T^{-\eta}$. For the BID we get an exponent of $\eta_{BID}=0.42$. For FCI and GRIDE we get much higher values, $\eta_{FCI}=0.89$, and $\eta_{GRIDE}=0.70$. Number of samples used: $N_{s}=5000$.
A.8.3 BID of MNIST and CIFAR10
In this section we compute the BID of images directly on data space, for the MNIST and CIFAR10 datasets. For MNIST, different ID estimatorsDenti et al. [2022], Spigler et al. [2020] predict values between 10 and 15. The natural binary representation of MNIST is obtained rounding the floating-point values of the gray scale to either black or white. We found that this operation increases the value of the intrinsic dimension when computed with standard approaches. For example, following Ref. Spigler et al. [2020], the intrinsic dimension (there called effective dimension, $d_{eff}$) is obtained as the scaling exponent relating the average nearest-neighbor distance $\langle\delta\rangle$ with the number of samples $N_{s}$, namely $\langle\delta\rangle\sim N_{s}^{-\frac{1}{d_{eff}}}$. Through a linear fit with $N_{s}\in[4000,8000]$, we get $d_{eff}(MNIST)\approx 16$ and $d_{eff}(round(MNIST))\approx 22$. The value obtained by our estimator is consistent with those numbers, as we found $BID(MNIST)\approx 19$.
In the case of colored images it is no longer trivial to find a meaningful binarization that approximately preserves the information. The rounding procedure used for MNIST applied on each channel of the RGB images of CIFAR10 gives a BID of the order of $200$, at odds with reported valuesPope et al. [2021], Denti et al. [2022], Spigler et al. [2020] ranging between 15 and 35. This discrepancy is not surprising since our estimator is specifically designed to work with binary systems, and by binarizing complex RGB images one losses a lot of information. In our main manuscript, together with Sec. A.8.1, we showed that the representations of natural images inside deep neural networks allow to be meaningfully binarized, at least with respect to the measurement of the intrinsic dimension. The construction of meaningful binary representations of colored images and their impact on the corresponding BID remains a very interesting open line of research.
A.9 Neighbourhoods are approximately preserved after binarizing high dimensional activations